OpenStudy (anonymous):

A small airplane traveling east, with the jet stream can cover 525 miles in 3 hours. Going west, it takes the plane 7 hours to go the same distance. How fast can the plane fly in still air?

OpenStudy (anonymous):

\[r=\frac{d}{t} \]Let r be the speed of the plane in still air and j be the speed of the jet stream. Solve the following equations for r and j:\[\left\{r+j=\frac{525}{3},(r-j)=\frac{525}{7}\right\}\]\[\{r=125\text{ mph},j=50\text{ mph}\} \]